IF A1,A2,A3,...,A BE N-TH ROOTS OF THE UNITY THEN THE VALUE OF (A1/3-A1+A2/3-A2+A3/3-A3+...+A/3-A)=


A)N/3N-1 


B)N-1/3N-1 


Grade:11

1 Answers

A.S. Rajavel
12 years ago

sum of roots is zero.

therefore -a1 -a2-a3......=0

1/3(0)=0

0+0=0